OpenStudy (anonymous):

SOMEONE PLEASE HELP ME. I WILL GIVE MEDAL AND BECOME FAN. A bottle of lotion is shaped like a cylinder with a radius of 3 cm and a height of 12 cm. The empty tube inside the cylinder has a radius of 0.5 cm and is the same height as the bottle. What is the volume of the lotion in the bottle? Use 3.14 to approximate pi, and express your answer to one decimal place.

OpenStudy (anonymous):

Well, first we need to find the volume of the outside cylinder: \[V = \pi r ^{2}h\] Which, in this equation, is: \[V = \pi (3)^{2}(12)\] Which, when put into a calculator, equals roughly 339. Next, we find the volume of the inside cylinder: \[V = \pi (0.5)^{2}(12)\] Which, when put into a calculator, equals: 9.42. So, I suppose the amount of lotion left inside the bottle would be 9.42? I'm a bit confused by what it's asked but that's how I worked it out.

OpenStudy (anonymous):

Ok thank you so much!